A COMMENT ON HÖNL-LONDON FACTORS

ANNIE HANSSON, Stockholm University, AlbaNova University Center, Department of Physics, SE-10691 Stockholm, Sweden; \underline{JAMES K. G. WATSON}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada K1A 0R6.

The Hönl-London factors^{a} are the factors dependent on the rotational quantum numbers in formulas for intensities of rovibronic transitions. A table of Hönl-London factors for singlet-singlet transitions of a diatomic molecule was given by Herzberg^{b}, and reproduced in other books. Standard definitions of transition moments and sum rules were proposed by Whiting and Nicholls^{c}, but there does not seem to be a published table of the corresponding individual Hönl-London factors. Here we show that for rotational transitions between levels of definite parity the values in Herzberg’s table should be multiplied by 2 for all perpendicular vibronic transitions, with an additional factor of 2 for Π−Σ and Σ−Π transitions. The results are consistent with the Whiting-Nicholls conventions.^{c}

aH. Hönl and F. London, Z. Physik 33, 803 (1925).

bG. Herzberg, Spectra of Diatomic Molecules, Van Nostrand (1950), p. 208.

cE. E. Whiting and R. W. Nicholls, Astrophys. J. Suppl. Ser. 27, 1 (1974); E. E. Whiting, A. Schadee, J. B. Tatum, J. T.

Hougen and R. W. Nicholls, J. Mol. Spectrosc. 80, 249 (1980).